OpenStudy (anonymous):

-2 (3 – m) + 15 = 6m -4(m -20)

OpenStudy (anonymous):

can somebody help me with this. -6 – 6m + 15 = 6m – 4m – 80 m + 15 = 10m – 80 11m = 15 – 80 11m = -65 (This is okay) Please check it